# Boundary feedback control for hyperbolic systems

### v2

@@ Abstract
We are interested in the feedback stabilization of general linear multi-dimensional first order hyperbolic systems in $\mathbb{R}^{d}$. Using a Lyapunov function with a suited weight function dpending on the system under consideration we show stabilization in $L^{2}$ for the studied system using a suitable feedback control. Therefore the controlability of the studied system is related to the feasibility of an associated linear matrix inequality. We show the applicability discussing the barotropic Euler equations.
@@ 1 Introduction
However, to the best of our knowledge the presented results are limited to the spatially one–dimensional case. Based on an example in metal forming processes, see [1, 35] we extended results to multi–dimensional hyperbolic balance laws. A specific system in two dimensions is discussed in [20] where a control problem for the shallow water equations is studied. There the authors take advantage of the structure of the system and show that the energy is non-increasing upon imposing suited boundary conditions. Just very recently a preprint was published where the boundary stabilization for two dimensional systems is studied using a different Lyapunov function, see [51]. A further recent result on multi–dimensional hyperbolic scalar conservation laws was presented in [47]. However, the goal of the mentioned paper is different from the one presented here. In [47] the author discusses scalar conservation laws in $\mathbb{R}^{d}$ whereas we are concerned with systems in a bounded domain. Moreover, we consider deviations from a reference state and thus study the linearized system, whereas in [47] a situation is studied when two fixed states at infinity are prescribed. It will be interesting and part of future research to study the relation between the works [20, 51, 47] and the results presented in this work.
@@ 2 Stabilization of Multi-Dimensional Linear Hyperbolic Balance Laws
[…]
@@
Let $\mathbf{w}(t,\mathbf{x})\in C^{1}\left((0,T),H^{s}(\Omega)\right)^{n}$, $s\geq 1+d/2,$ be a solution to the IBVP (2.1). The problem (2.1) is boundary diagonal and admits a feasible Lyapunov potential. Thus there exist $\mu(\mathbf{x})\in H^{s}(\Omega)$ such that
@@
holds true for some value $C_{A}\in\mathbb{R}_{>0}$. Define the Lyapunov function as follows
@@
The boundary condition for (2.1) is given by
@@
$\begin{split}&-\sum_{i=1}^{n}\int_{\Gamma_{i}^{-}}\lambda_{i}(\mathbf{x},\mathbf{n}(\mathbf{x}))\tilde{u}_{i}(t,\mathbf{x})^{2}\exp(\mu(\mathbf{x}))\,\textup{d}\mathbf{x}\leq\sum_{i=1}^{n}\int_{\Gamma_{i}^{+}}\lambda_{i}(\mathbf{x},\mathbf{n}(\mathbf{x}))v_{i}(t,\mathbf{x})^{2}\exp(\mu(\mathbf{x}))\,\textup{d}\mathbf{x},\end{split}$ (2.9)
with $\mathbf{v}=\mathbf{T}^{T}\mathbf{w}$ and (2.2). Then, it holds for the Lyapunov function that
$\frac{\textup{d}}{\textup{d}t}L(t)\leq-C_{L}L(t),\,C_{L}\in\mathbb{R}_{>0}.$
@@ 3.1 Estimating the Boundary Integral
$\mathcal{B}(t)=\int_{\partial\Omega}\mathcal{A}(t,\mathbf{x})\cdot\mathbf{n}(\mathbf{x})\,\textup{d}\mathbf{x}=\int_{\partial\Omega}\mathbf{v}^{T}(t,\mathbf{x})\mathbf{\Lambda}^{\ast}(\mathbf{x},\mathbf{n}(\mathbf{x}))\mathbf{v}(t,\mathbf{x})\exp(\mu(\mathbf{x}))\,\textup{d}\mathbf{x}.$
@@
$\mathcal{B}(t)$ $=\int_{\partial\Omega}\mathcal{A}(\mathbf{x})\cdot\mathbf{n}(\mathbf{x})\,\textup{d}\mathbf{x}=\int_{\partial\Omega}\mathbf{v}^{T}(t,\mathbf{x})\mathbf{\Lambda}^{\ast}(\mathbf{x},\mathbf{n}(\mathbf{x}))\mathbf{v}(t,\mathbf{x})\exp(\mu(\mathbf{x}))\,\textup{d}\mathbf{x}$
$=\sum_{i=1}^{n}\int_{\Gamma_{i}^{-}}\lambda_{i}(\mathbf{x},\mathbf{n}(\mathbf{x}))\exp(\mu(\mathbf{x}))\tilde{u}_{i}(t,\mathbf{x})^{2}\,\textup{d}\mathbf{x}+\sum_{i=1}^{n}\int_{\Gamma_{i}^{+}}\lambda_{i}(\mathbf{x},\mathbf{n}(\mathbf{x}))\exp(\mu(\mathbf{x}))v_{i}(t,\mathbf{x})^{2}\,\textup{d}\mathbf{x}$
@@
Due to $\mathbf{v}=\mathbf{T}^{T}\mathbf{w}$ we have $\mathbf{u}=\mathbf{T}\tilde{\mathbf{u}}$.
@@ 3.2 Estimating the Volume Integral
$\mathcal{I}(t)$ $:=\int_{\Omega}\mathbf{w}^{T}(t,\mathbf{x})\left[\sum_{k=1}^{d}\left(m_{k}(\mathbf{x})\mathbf{A}^{(k)}(\mathbf{x})+\frac{\partial}{\partial x_{k}}\mathbf{A}^{(k)}(\mathbf{x})\right)-2\mathbf{B}(\mathbf{x})\right]\mathbf{w}(t,\mathbf{x})\exp(\mu(\mathbf{x}))\,\textup{d}\mathbf{x}$
@@
If (3.1) is less or equal to zero there is nothing to do and it can simply be estimated by zero. Assume it is strict positive. In the case that $\mathbf{B}$ is not symmetric it can be replaced by its symmetric part $\mathbf{B}^{sym}=(\mathbf{B}+\mathbf{B}^{T})/2$. Since the $\mathbf{A}^{(k)}$ are symmetric this property is inherited by the derivatives of these matrices. Hence the following estimate can be established
$\mathbf{w}^{T}(t,\mathbf{x})\left(\sum_{k=1}^{d}\frac{\partial}{\partial x_{k}}\mathbf{A}^{(k)}(\mathbf{x})-2\mathbf{B}^{sym}(\mathbf{x})\right)\mathbf{w}(t,\mathbf{x})\leq C_{B}\|\mathbf{w}(t,\mathbf{x})\|_{2}^{2},\,C_{B}\in\mathbb{R}_{>0}.$
@@
$\mathcal{I}(t)$ $:=\int_{\Omega}\mathbf{w}^{T}(t,\mathbf{x})\left[\sum_{k=1}^{d}\left(m_{k}(\mathbf{x})\mathbf{A}^{(k)}(\mathbf{x})+\frac{\partial}{\partial x_{k}}\mathbf{A}^{(k)}(\mathbf{x})\right)-2\mathbf{B}(\mathbf{x})\right]\mathbf{w}(t,\mathbf{x})\exp(\mu(\mathbf{x}))\,\textup{d}\mathbf{x}$
@@
Note that by introducing the $\exp\left(\mu^{(\alpha)}\right)$ we want to highlight that in principal it is possible to choose suited Lyapunov potentials for every occurring multi-index $|\alpha|\leq s$. Once the Lyapunov function is introduced the strategy is the same and due to the similar structure inherited by the governing PDE (2.1) the estimates can be adopted from the previous proof. We exemplary want to give the Lyapunov function in the case $H^{1}$
@@
During the proof we encounter the term (3.1) and in particular treat the situation when it is strictly positive. It is further stated that under the given assumptions, namely (2.7), it can simply be estimated by zero in the case that it is less or equal to zero. However, there may be situations when the estimate of the volume integral may benefit from this term. In the case that the reminder (3.1) should be taken into account we introduce
@@
    $-\sum_{i=1}^{n}\tilde{u}_{i}(t)^{2}\int_{\Gamma_{i}^{-}}\lambda_{i}(\mathbf{x},\mathbf{n}(\mathbf{x}))e^{\mu(\mathbf{x})}\,\textup{d}\mathbf{x}\leq\sum_{i=1}^{n}\int_{\Gamma_{i}^{+}}v_{i}(t,\mathbf{x})^{2}\lambda_{i}(\mathbf{x},\mathbf{n}(\mathbf{x}))e^{\mu(\mathbf{x})}\,\textup{d}\mathbf{x}.$
@@
    $-\sum_{i=1}^{n}\int_{\Gamma^{-}_{i}}\tilde{u}(t,\mathbf{x})^{2}\lambda(\mathbf{x},\mathbf{n}(\mathbf{x}))e^{\mu(\mathbf{x})}\,\textup{d}\mathbf{x}\leq\sum_{i=1}^{n}\int_{\Gamma_{i}^{+}}v_{i}(t,\mathbf{x})^{2}\lambda_{i}(\mathbf{x},\mathbf{n}(\mathbf{x}))e^{\mu(\mathbf{x})}\,\textup{d}\mathbf{x}.$
@@
    Assume that both previous assumptions hold, i.e. ß$tilde{u}_{k}\equiv\tilde{u}(t),\,i=1,\dots,n$. Then we can give an explicit condition for the feedback control as
    $\tilde{u}(t)^{2}\leq-\left(\sum_{i=1}^{n}\int_{\Gamma^{-}_{i}}\lambda_{i}(\mathbf{x},\mathbf{n}(\mathbf{x}))e^{\mu(\mathbf{x})}\,\textup{d}\mathbf{x}\right)^{-1}\sum_{i=1}^{n}\int_{\Gamma_{i}^{+}}v_{i}(t,\mathbf{x})^{2}\lambda_{i}(\mathbf{x},\mathbf{n}(\mathbf{x}))e^{\mu(\mathbf{x})}\,\textup{d}\mathbf{x}.$ (3.4)
@@
    $\tilde{u}(t)=C\sqrt{-\left(\sum_{i=1}^{n}\int_{\Gamma^{-}_{i}}\lambda_{i}(\mathbf{x},\mathbf{n}(\mathbf{x}))e^{\mu(\mathbf{x})}\,\textup{d}\mathbf{x}\right)^{-1}\sum_{i=1}^{n}\int_{\Gamma_{i}^{+}}v_{i}(t,\mathbf{x})^{2}\lambda_{i}(\mathbf{x},\mathbf{n}(\mathbf{x}))e^{\mu(\mathbf{x})}\,\textup{d}\mathbf{x}}$ (3.5)
@@ 4.1 Barotropic Euler Equations
$\begin{split}\frac{\partial}{\partial t}\rho+\nabla_{x}\cdot\left(\rho\mathbf{v}\right)&=0,\\ \frac{\partial}{\partial t}(\rho v_{i})+\nabla_{x}\cdot\left(\rho v_{i}\mathbf{v}+p\mathbf{e}^{(i)}\right)&=0,\;i\in\{1,2\}.\end{split}$ (4.3)
The fluxes are given by
@@
$\begin{split}\frac{\partial}{\partial t}\rho+\mathbf{v}\cdot\nabla_{x}\rho+\rho\nabla_{x}\cdot\mathbf{v}&=0,\\ \frac{\partial}{\partial t}v_{i}+\mathbf{v}\cdot\nabla_{x}v_{i}+\frac{a^{2}}{\rho}\frac{\partial}{\partial x_{i}}\rho&=0,\;i\in\{1,2\}.\end{split}$ (4.11)
The Jacobians are given by
@@
When linearizing the system we simply fix the Jacobians at the desired reference state. We want to study the half-space problem for the linearized Euler equations in $\Omega:=\{x_{1}>0\}\times\mathbb{R}$. Thus we have the boundary $\partial\Omega=\{x_{1}=0\}$ with the outward pointing unit normal $\mathbf{n}=(-1,0)$. It is now possible to transform the system such that $\mathbf{A}^{(1)}$ is a diagonal matrix. We consider the reference state $\overline{\rho}\in\mathbb{R}_{>0},\overline{\mathbf{v}}\in\mathbb{R}^{2}\setminus\{\mathbf{0}\}$. The Jacobians of the transformed system read
$\overline{\mathbf{A}}^{(1)}=\begin{pmatrix}\overline{v}_{1}-a&0&0\\ 0&\overline{v}_{1}&0\\ 0&0&\overline{v}_{1}+a\end{pmatrix}\quad\text{and}\quad\overline{\mathbf{A}}^{(2)}=\begin{pmatrix}\overline{v}_{2}&\dfrac{\overline{\rho}}{2}&0\\ \dfrac{a^{2}}{\overline{\rho}}&\overline{v}_{2}&\dfrac{a^{2}}{\overline{\rho}}\\ 0&\dfrac{\overline{\rho}}{2}&\overline{v}_{2}\end{pmatrix}.$ (4.25)
The linearized system can be symmetrized by applying the following variable transformation $\mathbf{w}\to\mathbf{w}^{\ast}=(r,\mathbf{u})^{T}$, see also [26],
$\rho\to r:=\frac{a}{\overline{\rho}}\rho,\quad v_{1}\to u_{1}:=\frac{v_{1}}{\sqrt{2}},\quad v_{2}\to u_{2}:=\frac{a}{\overline{\rho}}v_{2}.$
We hence obtain
$\frac{\partial}{\partial t}\mathbf{w}^{\ast}+\overline{\mathbf{A}}^{(1)}\frac{\partial}{\partial x_{1}}\mathbf{w}^{\ast}+\overline{\mathbf{B}}^{(2)}\frac{\partial}{\partial x_{2}}\mathbf{w}^{\ast}=0$
with the Jacobian
$\overline{\mathbf{B}}^{(2)}=\begin{pmatrix}\overline{v}_{2}&\dfrac{a}{\sqrt{2}}&0\\ \dfrac{a}{\sqrt{2}}&\overline{v}_{2}&\dfrac{a}{\sqrt{2}}\\ 0&\dfrac{a}{\sqrt{2}}&\overline{v}_{2}\end{pmatrix}.$ (4.29)
We now discuss the LMI
$C_{L}\mathbf{Id}+m_{1}\overline{\mathbf{A}}^{(1)}+m_{2}\overline{\mathbf{B}}^{(2)}\leq 0$ (4.30)
@@
$\mathbf{A}(\overline{\mathbf{m}})$ $:=-\mathbf{Id}+\overline{m}_{1}\overline{\mathbf{A}}^{(1)}+\overline{m}_{2}\overline{\mathbf{B}}^{(2)}\geq 0,$
$\mathbf{A}(\overline{\mathbf{m}})$ $=\begin{pmatrix}-1+\overline{m}_{1}(\overline{v}_{1}-a)+\overline{m}_{2}\overline{v}_{2}&\overline{m}_{2}\dfrac{a}{\sqrt{2}}&0\\ \overline{m}_{2}\dfrac{a}{\sqrt{2}}&-1+\overline{m}_{1}\overline{v}_{1}+\overline{m}_{2}\overline{v}_{2}&\overline{m}_{2}\dfrac{a}{\sqrt{2}}\\ 0&\overline{m}_{2}\dfrac{a}{\sqrt{2}}&-1+\overline{m}_{1}(\overline{v}_{1}+a)+\overline{m}_{2}\overline{v}_{2}\end{pmatrix}.$
Note that we have
$\begin{cases}-1+\overline{m}_{1}(\overline{v}_{1}-a)+\overline{m}_{2}\overline{v}_{2}<-1+\overline{m}_{1}\overline{v}_{1}+\overline{m}_{2}\overline{v}_{2}<-1+\overline{m}_{1}(\overline{v}_{1}+a)+\overline{m}_{2}\overline{v}_{2},\;\overline{m}_{1}a>0,\\ -1+\overline{m}_{1}(\overline{v}_{1}-a)+\overline{m}_{2}\overline{v}_{2}>-1+\overline{m}_{1}\overline{v}_{1}+\overline{m}_{2}\overline{v}_{2}>-1+\overline{m}_{1}(\overline{v}_{1}+a)+\overline{m}_{2}\overline{v}_{2},\;\overline{m}_{1}a<0.\end{cases}$
Now we determine $\overline{\mathbf{m}}$ such that we have
$\begin{cases}-1+\overline{m}_{1}(\overline{v}_{1}-a)+\overline{m}_{2}\overline{v}_{2}=\sqrt{2}|\overline{m}_{2}a|>0,\;\overline{m}_{1}a>0,\\ -1+\overline{m}_{1}(\overline{v}_{1}+a)+\overline{m}_{2}\overline{v}_{2}=\sqrt{2}|\overline{m}_{2}a|>0,\;\overline{m}_{1}a<0.\end{cases}$
With this choice we have positive diagonal entries and a strongly diagonal dominant matrix and hence the matrix $\mathbf{A}(\overline{\mathbf{m}})$ is positive definite. Therefore $-\mathbf{A}(\overline{\mathbf{m}})$ is negative definite and with the choice $\mathbf{m}=-C_{L}\overline{\mathbf{m}}$ the LMI (4.30) holds and we further set

### v3

@@ Abstract
We are interested in the feedback stabilization of general linear multi-dimensional first order hyperbolic systems in $\mathbb{R}^{d}$. Using a Lyapunov function with a suited weight function depending on the system under consideration we show stabilization in $L^{2}$ for the studied system using a suitable feedback control. Therefore the controllability of the studied system is related to the feasibility of an associated linear matrix inequality. We show the applicability discussing the barotropic Euler equations.
@@ 1 Introduction
However, to the best of our knowledge the presented results are limited to the spatially one–dimensional case. Based on an example in metal forming processes, see [1, 34] we extended results to multi–dimensional hyperbolic balance laws. A specific system in two dimensions is discussed in [20] where a control problem for the shallow water equations is studied. There the authors take advantage of the structure of the system and show that the energy is non-increasing upon imposing suited boundary conditions. Just very recently a preprint was published where the boundary stabilization for two dimensional systems is studied using a different Lyapunov function, see [50] and a comparative study is given in [35]. A further recent result on multi–dimensional hyperbolic scalar conservation laws was presented in [46]. However, the goal of the mentioned paper is different from the one presented here. In [46] the author discusses scalar conservation laws in $\mathbb{R}^{d}$ whereas we are concerned with systems in a bounded domain. Moreover, we consider deviations from a reference state and thus study the linearized system, whereas in [46] a situation is studied when two fixed states at infinity are prescribed. It will be interesting and part of future research to study the relation between the works [20, 50, 46] and the results presented in this work.
@@ 2 Stabilization of Multi-Dimensional Linear Hyperbolic Balance Laws
$\overline{\mathbf{m}}:=\nabla\overline{\mu}(\mathbf{x})\quad\text{and}\quad\overline{\mathbf{A}}(\mathbf{m}):=-\mathbf{Id}+\sum_{k=1}^{d}\overline{m}_{k}\mathbf{A}^{(k)}\geq 0.$ (2.5)
@@
Let $\mathbf{w}(t,\mathbf{x})\in C^{1}\left((0,T),H^{s}(\Omega)\right)^{n}$, $s\geq 1+d/2,$ be a solution to the IBVP (2.1). Assume the problem (2.1) admits a feasible Lyapunov potential. Thus there exist $\mu(\mathbf{x})\in H^{s}(\Omega)$ such that
@@
holds true for some value $C_{A}\in\mathbb{R}_{>0}$ with $C_{A}>C_{B}$ and $C_{B}$ given as in (3.2). Define the Lyapunov function as follows
@@
Let the boundary condition for (2.1) be given by
@@
$\begin{split}&-\sum_{i=1}^{n}\int_{\Gamma_{i}^{-}}\lambda_{i}(\mathbf{x},\mathbf{n}(\mathbf{x}))\tilde{u}_{i}(t,\mathbf{x})^{2}\exp(\mu(\mathbf{x}))\,\textup{d}\mathbf{x}\leq\sum_{i=1}^{n}\int_{\Gamma_{i}^{+}}\lambda_{i}(\mathbf{x},\mathbf{n}(\mathbf{x}))q_{i}(t,\mathbf{x})^{2}\exp(\mu(\mathbf{x}))\,\textup{d}\mathbf{x},\end{split}$ (2.9)
with $\mathbf{q}=\mathbf{T}^{T}\mathbf{w}$ and (2.2).
Then, it holds for the Lyapunov function that for $0<C_{L}=C_{A}-C_{B}$
$\frac{\textup{d}}{\textup{d}t}L(t)\leq-C_{L}L(t).$
@@ 3.1 Estimating the Boundary Integral
$\mathcal{B}(t)=\int_{\partial\Omega}\mathcal{A}(t,\mathbf{x})\cdot\mathbf{n}(\mathbf{x})\,\textup{d}\mathbf{x}=\int_{\partial\Omega}\mathbf{q}^{T}(t,\mathbf{x})\mathbf{\Lambda}^{\ast}(\mathbf{x},\mathbf{n}(\mathbf{x}))\mathbf{q}(t,\mathbf{x})\exp(\mu(\mathbf{x}))\,\textup{d}\mathbf{x}.$
@@
$\mathcal{B}(t)$ $=\int_{\partial\Omega}\mathcal{A}(\mathbf{x})\cdot\mathbf{n}(\mathbf{x})\,\textup{d}\mathbf{x}=\int_{\partial\Omega}\mathbf{q}^{T}(t,\mathbf{x})\mathbf{\Lambda}^{\ast}(\mathbf{x},\mathbf{n}(\mathbf{x}))\mathbf{q}(t,\mathbf{x})\exp(\mu(\mathbf{x}))\,\textup{d}\mathbf{x}$
$=\sum_{i=1}^{n}\int_{\Gamma_{i}^{-}}\lambda_{i}(\mathbf{x},\mathbf{n}(\mathbf{x}))\exp(\mu(\mathbf{x}))\tilde{u}_{i}(t,\mathbf{x})^{2}\,\textup{d}\mathbf{x}+\sum_{i=1}^{n}\int_{\Gamma_{i}^{+}}\lambda_{i}(\mathbf{x},\mathbf{n}(\mathbf{x}))\exp(\mu(\mathbf{x}))q_{i}(t,\mathbf{x})^{2}\,\textup{d}\mathbf{x}$
@@
Due to $\mathbf{q}=\mathbf{T}^{T}\mathbf{w}$ we have $\mathbf{u}=\mathbf{T}\tilde{\mathbf{u}}$.
@@ 3.2 Estimating the Volume Integral
$\mathcal{I}(t)$ $=\int_{\Omega}\mathbf{w}^{T}(t,\mathbf{x})\left[\sum_{k=1}^{d}\left(m_{k}(\mathbf{x})\mathbf{A}^{(k)}(\mathbf{x})+\frac{\partial}{\partial x_{k}}\mathbf{A}^{(k)}(\mathbf{x})\right)-2\mathbf{B}(\mathbf{x})\right]\mathbf{w}(t,\mathbf{x})\exp(\mu(\mathbf{x}))\,\textup{d}\mathbf{x}$
@@
If (3.1) is less or equal to zero there is nothing to do and it can simply be estimated by zero. Assume (3.1) is strictly positive. In the case that $\mathbf{B}$ is not symmetric it can be replaced by its symmetric part $\mathbf{B}^{sym}=(\mathbf{B}+\mathbf{B}^{T})/2$. Since the $\mathbf{A}^{(k)}$ are symmetric this property is inherited by the derivatives of these matrices. Hence the following estimate can be established
$\mathbf{w}^{T}(t,\mathbf{x})\left(\sum_{k=1}^{d}\frac{\partial}{\partial x_{k}}\mathbf{A}^{(k)}(\mathbf{x})-2\mathbf{B}^{sym}(\mathbf{x})\right)\mathbf{w}(t,\mathbf{x})\leq C_{B}\|\mathbf{w}(t,\mathbf{x})\|_{2}^{2},\,C_{B}\in\mathbb{R}_{>0}.$ (3.2)
@@
$\mathcal{I}(t)$ $=\int_{\Omega}\mathbf{w}^{T}(t,\mathbf{x})\left[\sum_{k=1}^{d}\left(m_{k}(\mathbf{x})\mathbf{A}^{(k)}(\mathbf{x})+\frac{\partial}{\partial x_{k}}\mathbf{A}^{(k)}(\mathbf{x})\right)-2\mathbf{B}(\mathbf{x})\right]\mathbf{w}(t,\mathbf{x})\exp(\mu(\mathbf{x}))\,\textup{d}\mathbf{x}$
@@
Note that by introducing the $\exp\left(\mu^{(\alpha)}\right)$ we want to highlight that in principle it is possible to choose suited Lyapunov potentials for every occurring multi-index $|\alpha|\leq s$. Each Lyapunov potential has to satisfy a LMI similar to (2.3) with respect to the system corresponding to the studied multi-index. It has to be studied under which conditions this is possible for a given system. Once the Lyapunov function is introduced the strategy is the same and due to the similar structure inherited by the governing PDE (2.1) the estimates can be adopted from the previous proof. We exemplary want to give the Lyapunov function in the case $H^{1}$
@@
During the proof we encounter the term (3.1). In particular in the proof we only treat the situation when this expression is strictly positive since under the given assumptions, namely (2.7), it can simply be estimated by zero in the case that it is less or equal to zero. However, there may be situations when the estimate of the volume integral may benefit from this term. In the case that the reminder (3.1) should be taken into account we introduce
@@
    $-\sum_{i=1}^{n}\tilde{u}_{i}(t)^{2}\int_{\Gamma_{i}^{-}}\lambda_{i}(\mathbf{x},\mathbf{n}(\mathbf{x}))e^{\mu(\mathbf{x})}\,\textup{d}\mathbf{x}\leq\sum_{i=1}^{n}\int_{\Gamma_{i}^{+}}q_{i}(t,\mathbf{x})^{2}\lambda_{i}(\mathbf{x},\mathbf{n}(\mathbf{x}))e^{\mu(\mathbf{x})}\,\textup{d}\mathbf{x}.$
@@
    $-\sum_{i=1}^{n}\int_{\Gamma^{-}_{i}}\tilde{u}(t,\mathbf{x})^{2}\lambda_{i}(\mathbf{x},\mathbf{n}(\mathbf{x}))e^{\mu(\mathbf{x})}\,\textup{d}\mathbf{x}\leq\sum_{i=1}^{n}\int_{\Gamma_{i}^{+}}q_{i}(t,\mathbf{x})^{2}\lambda_{i}(\mathbf{x},\mathbf{n}(\mathbf{x}))e^{\mu(\mathbf{x})}\,\textup{d}\mathbf{x}.$
@@
    Assume that both previous assumptions hold, i.e. $\tilde{u}_{i}\equiv\tilde{u}(t),\,i=1,\dots,n$. Then we can give an explicit condition for the feedback control as
    $\tilde{u}(t)^{2}\leq-\left(\sum_{i=1}^{n}\int_{\Gamma^{-}_{i}}\lambda_{i}(\mathbf{x},\mathbf{n}(\mathbf{x}))e^{\mu(\mathbf{x})}\,\textup{d}\mathbf{x}\right)^{-1}\sum_{i=1}^{n}\int_{\Gamma_{i}^{+}}q_{i}(t,\mathbf{x})^{2}\lambda_{i}(\mathbf{x},\mathbf{n}(\mathbf{x}))e^{\mu(\mathbf{x})}\,\textup{d}\mathbf{x}.$ (3.5)
@@
    $\tilde{u}(t)=C\sqrt{-\left(\sum_{i=1}^{n}\int_{\Gamma^{-}_{i}}\lambda_{i}(\mathbf{x},\mathbf{n}(\mathbf{x}))e^{\mu(\mathbf{x})}\,\textup{d}\mathbf{x}\right)^{-1}\sum_{i=1}^{n}\int_{\Gamma_{i}^{+}}q_{i}(t,\mathbf{x})^{2}\lambda_{i}(\mathbf{x},\mathbf{n}(\mathbf{x}))e^{\mu(\mathbf{x})}\,\textup{d}\mathbf{x}}$ (3.6)
@@ 4.1 Barotropic Euler Equations
$\begin{split}\frac{\partial}{\partial t}\rho+\nabla_{x}\cdot\left(\rho\mathbf{v}\right)&=0,\\ \frac{\partial}{\partial t}(\rho v_{i})+\nabla_{x}\cdot\left(\rho v_{i}\mathbf{v}+p\mathbf{e}^{(i)}\right)&=0,\;i\in\{1,2\}\end{split}$ (4.3)
where $\rho>0$ denotes the mass density, $\mathbf{v}=(v_{1},v_{2})^{T}$ is the velocity and $p$ is the pressure as function of the density. The fluxes are given by
@@
$\begin{split}\frac{\partial}{\partial t}\rho+\mathbf{v}\cdot\nabla_{x}\rho+\rho\nabla_{x}\cdot\mathbf{v}&=0,\\ \frac{\partial}{\partial t}v_{i}+\mathbf{v}\cdot\nabla_{x}v_{i}+\frac{a^{2}}{\rho}\frac{\partial}{\partial x_{i}}\rho&=0,\;i\in\{1,2\}\end{split}$ (4.5)
with $a>0$ denoting the speed of sound. The Jacobians are given by
@@
When linearizing the system we simply fix the Jacobians at the desired reference state. We consider the reference state $\overline{\rho}\in\mathbb{R}_{>0},\overline{\mathbf{v}}\in\mathbb{R}^{2}\setminus\{\mathbf{0}\}$. The linearized system can be symmetrized by applying the following variable transformation $\mathbf{w}\to\mathbf{w}^{\ast}=(r,\mathbf{v})^{T}$
$\rho\to r:=\frac{\overline{a}}{\overline{\rho}}\rho,$
see also [25]. We hence obtain
$\frac{\partial}{\partial t}\mathbf{w}^{\ast}+\overline{\mathbf{A}}^{(1)}\frac{\partial}{\partial x_{1}}\mathbf{w}^{\ast}+\overline{\mathbf{A}}^{(2)}\frac{\partial}{\partial x_{2}}\mathbf{w}^{\ast}=0$
with the Jacobians
$\overline{\mathbf{A}}^{(1)}=\begin{pmatrix}\overline{v}_{1}&\overline{a}&0\\ \overline{a}&\overline{v}_{1}&0\\ 0&0&\overline{v}_{1}\end{pmatrix}\quad\text{and}\quad\overline{\mathbf{A}}^{(2)}=\begin{pmatrix}\overline{v}_{2}&0&\overline{a}\\ 0&\overline{v}_{2}&0\\ \overline{a}&0&\overline{v}_{2}\end{pmatrix}.$ (4.7)
Since the Jacobians are constant and $\mathbf{B}\equiv\mathbf{0}$ we do not have to discuss the term (3.2). Thus we have $C_{B}=0$ and set $C_{L}:=C_{A}>0$. We now discuss the LMI
$C_{L}\mathbf{Id}+m_{1}\overline{\mathbf{A}}^{(1)}+m_{2}\overline{\mathbf{A}}^{(2)}\leq 0$ (4.8)
@@
$\mathbf{A}(\overline{\mathbf{m}})$ $:=-\mathbf{Id}+\overline{m}_{1}\overline{\mathbf{A}}^{(1)}+\overline{m}_{2}\overline{\mathbf{A}}^{(2)}\geq 0,$
$\mathbf{A}(\overline{\mathbf{m}})$ $=\begin{pmatrix}-1+\overline{m}_{1}\overline{v}_{1}+\overline{m}_{2}\overline{v}_{2}&\overline{m}_{1}\overline{a}&\overline{m}_{2}\overline{a}\\ \overline{m}_{1}\overline{a}&-1+\overline{m}_{1}\overline{v}_{1}+\overline{m}_{2}\overline{v}_{2}&0\\ \overline{m}_{2}\overline{a}&0&-1+\overline{m}_{1}\overline{v}_{1}+\overline{m}_{2}\overline{v}_{2}\end{pmatrix}.$
A sufficient criterion would be to choose $\overline{m}_{1}$ and $\overline{m}_{2}$ such that $\mathbf{A}(\overline{\mathbf{m}})$ becomes strongly diagonal dominant, i.e., the diagonal entries are strict positive and larger than the sum of the absolute row entries. Due to the special structure of the matrix we hence look for coefficients with the property that
$-1+\overline{m}_{1}\overline{v}_{1}+\overline{m}_{2}\overline{v}_{2}=\sigma\left(|\overline{m}_{1}|+|\overline{m}_{2}|\right)\overline{a},\;\sigma\in(1,\infty).$
With this choice we have
$-1+\overline{m}_{1}\overline{v}_{1}+\overline{m}_{2}\overline{v}_{2}>\left(|\overline{m}_{1}|+|\overline{m}_{2}|\right)\overline{a}\geq\max\{|\overline{m}_{1}|,|\overline{m}_{2}|\}\overline{a}\geq\min\{|\overline{m}_{1}|,|\overline{m}_{2}|\}\overline{a}\geq 0$
and hence the matrix $\mathbf{A}(\overline{\mathbf{m}})$ is positive definite. Therefore $-\mathbf{A}(\overline{\mathbf{m}})$ is negative definite and with the choice $\mathbf{m}=-C_{L}\overline{\mathbf{m}}$ the LMI (4.8) holds and we further set